OpenStudy (anonymous):

What are the steps to simplifying the square root of 18?

OpenStudy (anonymous):

First rewrite: \[\sqrt{18}=\sqrt{9*2}=\sqrt{9}*\sqrt{2}=3*\sqrt{2}=3\sqrt2\]